How do you answer this question: the sum of 3 numbers is 79. The second number is 9 times the first, and the third number is 3 more than the second. Find the numbers.
Answer by matineesuxxx(27)

x + y + z = 79
the trick is writing both 'x' and 'z' in terms of 'y' so we only have one variable.
"the second number is 9 times the first" ---> y = 9x , therefore, x = y/9
"the third number is three more than the second" ----> z = 3+ y
now that we have written each variable in terms of "Y", you can plug them into our equation, 'x+y+z=79' and we get;
(y/9) + y + (3 + y) = 79
(y/9) + 2y + 3 = 79
now find a common denominator by multiplying both (2y/1) and (3/1) by (9/9)
(y+18y+27)/9=27
19y + 27 = 711
19y = 684
y = 36

therefore x = (36/9) = 4
z = 3 + 36 = 39

PROOF: 36 + 39 + 4 = 79